A 10×10^-6F capacitor is charged to a pd of 100V. A 2×10^-6F capacitor is connected in parallel with it, disconnected and discharged. This process of connection, disconnection and discharge is carried out four times. Calculate :

  1. The fraction of the original charge which remains on the larger capacitor after the first connection.
  2. The fraction of the original charge that remains after the fourth connection.
  3. The final pd across the 10×10^-6 capacitor
1
Expert's answer
2022-01-31T10:44:32-0500

"Part\>(i)\\

Q=CV"


In parallel arrangement, V is constant

"\implies\>Q\propto\>C"


Total capacitance in parallel "=10+2"

"=12\mu\>F"



Fraction of original charge on larger capacitor

"=\frac{10}{12}=\frac{5}{6}"


Part(ii)


(ii) 2nd connection:

Total charge "=(10×100)+\frac{1}{6}×1000"

"=\frac{7}{6}×1000"


(iii) 3rd Connection


Total charge "=1000+\frac{1}{6}(\frac{7}{6}×1000)"

"=\frac{43}{36}×1000"



4th Connection


Total charge "=1000+\frac{1}{6}×(\frac{43}{36}×1000)"

"=\frac{259}{216}×1000"


Fraction of the original charge that remains after the fourth connection


On small capacitor "=\frac{1}{6}×\frac{259}{216}"

"=\frac{259}{1296}"



On large capacitor "=\frac{5}{6}×\frac{259}{216}"

"=\frac{1295}{1296}"


Part(iii)



Final p.d "=\frac{Q}{C}=\frac{\frac{1295}{1296}×100}{10}"

"=99.92V"